OpenStudy (anonymous):

roots for a quadratic equation using the quadratic formula:

OpenStudy (anonymous):

\[32x ^{2}+100x+63\]

OpenStudy (anonymous):

Must be equal to 0..

OpenStudy (anonymous):

formula :\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

o yes sorry forgot the zero

OpenStudy (anonymous):

Yes.. Can you use this formula??

OpenStudy (anonymous):

what they said factor.. not find the roots

OpenStudy (anonymous):

can i still use the formula?

OpenStudy (anonymous):

But read your question..

OpenStudy (anonymous):

Question is to find the roots..

OpenStudy (anonymous):

i made a mistake, the book said factor the give trinomial

OpenStudy (anonymous):

Normal factorization becomes lengthy I guess..

OpenStudy (anonymous):

yes but using the formula would be wrong right? because i used the formula and got it wrong until I did normal factorization..

OpenStudy (anonymous):

\(32*63 = 2016\) Can you make its factors such that sum will be 100..

OpenStudy (anonymous):

okay, i did that but why did you multiply the coefficients?

OpenStudy (anonymous):

See in case: \[ax^2 + bx + c = 0\] If a is 1, then we simply make factors of c.. If a is not 1 then we should make factors of ac..

OpenStudy (anonymous):

oh yes: ____ x____ = ac __ + ____= b

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

Got the factors or not? I have got.. But I will let you try once..

OpenStudy (anonymous):

okay I got the answer using another method: factors of 32: (1,32) (4,8) (2,16) factors of 63: (1,63) (9,7) etc then multiply the factors: (4)(7) + (9)(8) = b = 100 then xs together and ys together (4x+9)(8x+7) =)

OpenStudy (anonymous):

I got the factors as 28 and 72.. Yes you got the same..

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

Well Done you are right..

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

thank you, I think what confused me was the factoring and the roots.. because when I used the formula it came out wrong

OpenStudy (anonymous):

Using quadratic formula ??

OpenStudy (anonymous):

yes, maybe I went wrong somewhere.. but if I used the formula I was looking for roots and not factors

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

okay (:

OpenStudy (anonymous):

\[D = b^2 - 4ac = 1936\] \[\sqrt{D} = \sqrt{1936} = 44\] \[x = \frac{-100 \pm 44}{2 \cdot 32}\] \[\implies x = -\frac{9}{4}\] \[\implies x = -\frac{7}{8}\] Yes it is coming right..

OpenStudy (anonymous):

okay you are write, I was getting weird decimals thats why.. so either method works :)

OpenStudy (anonymous):

They are both same.. This will never happen that you will get different answers for both.. Remember this..